# Entropy-Regularized Token-Level Policy Optimization for Large Language Models

### 09 Feb 2024 v1

@@ Abstract
Large Language Models (LLMs) have shown promise as intelligent agents in interactive decision-making tasks. Traditional approaches often depend on meticulously designed prompts, high-quality examples, or additional reward models for in-context learning, supervised fine-tuning, or RLHF. Reinforcement learning (RL) presents a dynamic alternative for LLMs to overcome these dependencies by engaging directly with task-specific environments. Nonetheless, it faces significant hurdles: 1) instability stemming from the exponentially vast action space requiring exploration; 2) challenges in assigning token-level credit based on action-level reward signals, resulting in discord between maximizing rewards and accurately modeling corpus data. In response to these challenges, we introduce Entropy-Regularized Token-level Policy Optimization (ETPO), an entropy-augmented RL method tailored for optimizing LLMs at the token level. At the heart of ETPO is our novel per-token soft Bellman update, designed to harmonize the RL process with the principles of language modeling. This methodology decomposes the Q-function update from a coarse action-level view to a more granular token-level perspective, backed by theoretical proof of optimization consistency. Crucially, this decomposition renders linear time complexity in action exploration. We assess the effectiveness of ETPO within a simulated environment that models data science code generation as a series of multi-step interactive tasks; results show that ETPO achieves effective performance improvement on the CodeLlama-7B model and surpasses a variant PPO baseline inherited from RLHF. This underlines ETPO’s potential as a robust method for refining the interactive decision-making capabilities of LLMs.
@@ 1 Introduction
The capability of large language models (LLMs) to function as intelligent agents has been showcased across various domains, such as mathematical research (Romera-Paredes et al., 2023), data science (Hollmann et al., 2023), software engineering (Qian et al., 2023), and geoscience (Deng et al., 2023; Lin et al., 2023). Given these developments, there is an increasing interest in enhancing their functionality in interactive decision-making contexts (Shridhar et al., 2020). Historically, the enhancement of LLMs has relied heavily on the crafting of precise prompts and the provision of high-quality examples for in-context learning or supervised fine-tuning (SFT) (Yao et al., 2022; Shinn et al., 2023; Radford et al., 2018) Fortunately, employing reinforcement learning (RL) (Sutton & Barto, 2018) offers a dynamic approach for LLMs to actively participate and progressively refine their performance in task-specific environments through a process of exploration and feedback. This shift towards RL allows for a reduction in dependency on manually prepared examples and prompts, paving the way for more autonomous and adaptive learning capabilities in LLMs.
@@ 2.2 Entropy-Regularized Reinforcement Learning
Entropy-regularized reinforcement learning (ERL) methods (Schulman et al., 2017a; Ziebart, 2010; Fox et al., 2015; Haarnoja et al., 2017; Nachum et al., 2017) are a subset of reinforcement learning that aims to search for the optimal stochastic policy under the constraints of a given distribution. ERL constrains the learned policy from deviating too far from a reference distribution by adding an entropy-regularized term to the reward function and optimizing with soft Bellman update (Schulman et al., 2017a). A concrete instance of ERL is the Soft Actor-Critic (Haarnoja et al., 2018a, b) algorithm, which defines the $\bar{\pi}$ to be a uniform distribution to balance the trade-off between exploration (entropy maximization) and exploitation (reward maximization). While PPO and some other Actor-Critic methods (Konda & Tsitsiklis, 1999) also encourage exploration by adding an entropy term as a penalty to the loss function, this penalty only takes effect in the current time step. RLHF (Ziegler et al., 2019) that adopts a variant of PPO with a KL term added to its reward function could also be seen as an approximation to ERL.
@@ 3.2 Entropy-Regularized Reinforcement Learning
$\mathcal{G}^{\prime}(\pi)=\mathbb{E}_{a\sim\pi}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}(r(s_{t},a_{t})\Big{]},$ (1)
@@
$\mathcal{G}(\pi)=\mathbb{E}_{a\sim\pi}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}(r(s_{t},a_{t})-\beta D_{KL}[\pi||\bar{\pi}](s_{t})\Big{]},$ (2)
@@ 4 Entropy-Regularized Token-Level Policy Optimization
$Q(s_{t},w_{t}^{1:j-1},w_{t}^{j})=\begin{cases}\mathbb{E}_{w_{t}^{j+1}\sim\pi(s_{t},w_{t}^{1:j})}[Q(s_{t},w_{t}^{1:j},w_{t}^{j+1})]-\beta D_{KL}[\pi||\bar{\pi}](s_{t},w_{t}^{1:j}),&\mbox{if }j<|a_{t}|\\ r(s_{t},a_{t})+\gamma(\mathbb{E}_{w_{t+1}^{1}\sim\pi(s_{t+1})}[Q(s_{t+1},w_{t+1}^{1})]-\beta D_{KL}[\pi||\bar{\pi}](s_{t+1}),&\mbox{if }j=|a_{t}|\end{cases}$ (6)
@@ 5.2 Main Results
Figure 3 shows the trends of average performance of ETPO and three related baseline methods across $14$ datasets under different exploration intensities, i.e. $k$. More specifically, Table 2 shows the highest reward explored by these methods for each dataset within $500$ interactive steps. Among these baselines, the Reflection keeps the interactive pipeline of ETPO unchanged but removes the training process as a reference for LLMs’ initial performance. Besides, similar to RLHF, we also implement an action-level PPO with KL term in the reward function: $r(s_{t},a_{t})-\beta\log\frac{\pi(a_{t}|s_{t})}{\rho(a_{t}|s_{t})}$, based on the same framework as ETPO to fine-tune CodeLlama-7B model, i.e. (PPO-KL). As an action-level algorithm, PPO-KL will assign the same credit to all tokens in an action.
@@ 5.5 Affect on Perplexity
Table 3: Perplexity changes comparison.
According to the experimental results (as shown in Table 3), the minor changes observed after training the CodeLlama-7B model using the ETPO method indicate that this method has a relatively small impact on the original language modeling. This slight variation suggests that although the ETPO method may introduce new optimization or training strategies, it does not fundamentally alter or disrupt the basic logic of language modeling. This is a positive outcome, as it means that while trying to improve model performance or introduce new training methods, the model’s fundamental capabilities and stability are maintained.

### 05 Mar 2024 v2

@@ Abstract
Large Language Models (LLMs) have shown promise as intelligent agents in interactive decision-making tasks. Recent approaches often depend on meticulously designed prompts, high-quality examples, or additional reward models for in-context learning, supervised fine-tuning, or RLHF. Reinforcement learning (RL) presents a dynamic alternative for LLMs to overcome these dependencies by engaging directly with task-specific environments. Nonetheless, it faces significant hurdles: 1) instability stemming from the exponentially vast action space requiring exploration; 2) challenges in assigning token-level credit based on action-level reward signals, resulting in discord between maximizing rewards and accurately modeling corpus data. In response to these challenges, we introduce Entropy-Regularized Token-level Policy Optimization (ETPO), an entropy-augmented RL method tailored for optimizing LLMs at the token level. At the heart of ETPO is our novel per-token soft Bellman update, harmonizing RL process with the principles of language modeling. It decomposes the Q-function update from a coarse action-level view to a more granular token-level perspective, backed by theoretical proof of optimization consistency. Crucially, this decomposition renders linear time complexity in action exploration. We assess the effectiveness of ETPO within a simulated environment that models data science code generation as a series of multi-step interactive tasks; results show that ETPO achieves effective performance improvement on the CodeLlama-7B model and surpasses a variant PPO baseline inherited from RLHF. This underlines ETPO’s potential as a robust method for refining the interactive decision-making capabilities of LLMs. Our code is open-sourced at https://github.com/morning9393/ETPO.
@@ 1 Introduction
The capability of large language models (LLMs) to function as intelligent agents has been showcased across various domains, such as mathematical research (Romera-Paredes et al., 2023), data science (Hollmann et al., 2023), software engineering (Qian et al., 2023), and geoscience (Deng et al., 2023; Lin et al., 2023). Given these developments, there is an increasing interest in enhancing their functionality in interactive decision-making contexts (Shridhar et al., 2020). Historically, the enhancement of LLMs has relied heavily on the crafting of precise prompts and the provision of high-quality examples for in-context learning or supervised fine-tuning (SFT) (Yao et al., 2022; Shinn et al., 2023; Radford et al., 2018). Fortunately, employing reinforcement learning (RL) (Sutton & Barto, 2018) offers a dynamic approach for LLMs to actively participate and progressively refine their performance in task-specific environments through a process of exploration and feedback. This shift towards RL allows for a reduction in dependency on manually prepared examples and prompts, paving the way for more autonomous and adaptive learning capabilities in LLMs.
@@ 2.2 Entropy-Regularized Reinforcement Learning
j methods (Schulman et al., 2017a; Ziebart, 2010; Fox et al., 2015; Haarnoja et al., 2017; Nachum et al., 2017) are a subset of reinforcement learning that aims to search for the optimal stochastic policy under the constraints of a given distribution. ERL constrains the learned policy from deviating too far from a reference distribution by adding an entropy-regularized term to the reward function and optimizing with soft Bellman update (Schulman et al., 2017a). A concrete instance of ERL is the Soft Actor-Critic (Haarnoja et al., 2018a, b) algorithm, which defines the $\bar{\pi}$ to be a uniform distribution to balance the trade-off between exploration (entropy maximization) and exploitation (reward maximization). While PPO and some other Actor-Critic methods (Konda & Tsitsiklis, 1999) also encourage exploration by adding an entropy term as a penalty to the loss function, this penalty only takes effect in the current time step. RLHF (Ziegler et al., 2019) that adopts a variant of PPO with a KL term added to its reward function could also be seen as an approximation to ERL.
@@ 3.2 Entropy-Regularized Reinforcement Learning
$\mathcal{G}^{\prime}(\pi)=\mathbb{E}_{a\sim\pi}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\Big{]},$ (1)
@@
$\mathcal{G}(\pi)=\mathbb{E}_{a\sim\pi}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}(r(s_{t},a_{t})-\beta D_{KL}[\pi||\bar{\pi}](s_{t}))\Big{]},$ (2)
@@ 4 Entropy-Regularized Token-Level Policy Optimization
$Q(s_{t},w_{t}^{1:j-1},w_{t}^{j})=\begin{cases}\mathbb{E}_{w_{t}^{j+1}\sim\pi(s_{t},w_{t}^{1:j})}[Q(s_{t},w_{t}^{1:j},w_{t}^{j+1})]-\beta D_{KL}[\pi||\bar{\pi}](s_{t},w_{t}^{1:j}),&\mbox{if }j<|a_{t}|\\ r(s_{t},a_{t})+\gamma(\mathbb{E}_{w_{t+1}^{1}\sim\pi(s_{t+1})}[Q(s_{t+1},w_{t+1}^{1})]-\beta D_{KL}[\pi||\bar{\pi}](s_{t+1})),&\mbox{if }j=|a_{t}|\end{cases}$ (6)
@@ 5.2 Main Results
Figure 3 shows the trends of average performance of ETPO and three related baseline methods across $14$ datasets under different exploration intensities, i.e. $k$. More specifically, Table 2 shows the highest reward explored by these methods for each dataset within $500$ interactive steps. Among these baselines, the Reflection keeps the interactive pipeline of ETPO unchanged but removes the training process as a reference for LLMs’ initial performance. Besides, similar to RLHF, we also implement an action-level PPO with KL term in the reward function: $r(s_{t},a_{t})-\beta\log\frac{\pi(a_{t}|s_{t})}{\rho(a_{t}|s_{t})}$, based on the same framework as ETPO to fine-tune CodeLlama-7B model, i.e. PPO-KL. As an action-level algorithm, PPO-KL will assign the same credit to all tokens in an action.
@@ 5.5 Affect on Perplexity
According to the experimental results in Table 3, the minor changes observed after training the CodeLlama-7B model using the ETPO method indicate that this method has a relatively small impact on the original language modeling. This slight variation suggests that although the ETPO method may introduce new optimization or training strategies, it does not fundamentally alter or disrupt the basic logic of language modeling. This is a positive outcome, as it means that while trying to improve model performance or introduce new training methods, the model’s fundamental capabilities and stability are maintained.